Find the domain of the rational function, then find the values of the function if possible.
T(x)=(5-x)/(x-5),T(5),T(-9)
I have answered as The domain is (-∞<5<∞) and the values of the function T(5)is 0 and undefined and the value of the function T(-9) is (-1,∞).
Is this right?

Answer by stanbon(75887)

Find the domain of the rational function, then find the values of the function if possible.
T(x)=(5-x)/(x-5) = -1(x-5)/(x-5) = -1
-------------
Since x cannot be 5, Domain is "All Real Numbers except 5"; (-oo, 5)U(5,+oo)
-------------------
T(5) is undefined
-----------------
T(-9) = (5+9)/(-9-5) = 14/(-14) = -1
--------------
Cheers,
Stan H.
-----------